# Neutrino Interaction Vertex Reconstruction in DUNE with Pandora Deep Learning

### v1

@@ 1 Introduction
Section 2 briefly introduces LArTPCs, while Section 3 describes the simulated data used for this work. The conceptual approach to finding the vertex and the corresponding truth definition are presented in Section 4. Section 5 summarises the network architecture, provides details of the training metrics and assesses the behaviour of the loss function and network weights. Results in a simulated experimental environment are presented in Section 6. The performance characteristics, sources of bias and model dependence considerations are discussed in Section 7 and a number of extensions/alternative approaches planned for future work are described in Section 8.
@@ 2 Liquid argon time projection chambers
Two far detector designs currently exist for DUNE, the horizontal-drift [7] and vertical-drift [8] detectors. These two designs utilise different geometries, orientations, drift lengths and readout technologies, the details of which are left to the aforementioned references. Conceptually the operation is the same, and Pandora is agnostic to the detector differences. A LArTPC volume is a fully active liquid argon target, with a uniform electric field applied between a cathode and anode. Charged particles propagating through this medium ionise the argon (see Fig. 1), producing drift electrons and scintillation photons. The drift electrons are carried by the electric field to the anode where they are read out on three readout planes: Two planes (U and V) with voltage biased to be transparent to the electrons and thereby allow current to be induced in readout channels as the electrons pass by, and a readout plane (W) with voltage set to permit collection of the drift electrons on the readout channels. The scintillation light is detected by photon detectors embedded in anode and cathode planes [9], though Pandora does not currently make use of scintillation light and therefore this will not be discussed further in this article.
@@ 5 Training and understanding the network
Given the previously described formulation of the problem in terms of semantic segmentation, the architecture used has a dense U-ResNet structure [19, 21]. The network has four blocks in each of the down-sampling and up-sampling paths, with all convolutional filters of size 3x3. The convolutional blocks implement residual skip connections between the input to those blocks and the final ReLU activation function of the respective block. Group normalization is applied within convolutional and transpose convolutional blocks rather than batch normalization due to the limited training batch size of 32 images per batch required to fit within GPU memory (16 GB) during training.
@@
For the second pass training, with equivalent inclusion cuts applied, there were 85981, 85555, and 84658 for the U, V and W views, respectively. Note that images are excluded if errors in the first pass produce framing of the candidate vertex region that does not contain the true vertex. For validation there were 28660, 28518, and 28219 for the U, V and W views, respectively. Each network was trained for 25 epochs and the evolution of the network performance across each of those training iterations is shown in Fig. 4(a)-4(d).
@@
Also of interest are the confusion matrices for the various classes of the different networks. As already noted, per class accuracy is high, but one can also observe the distribution of classification errors in Fig. 5(a)-5(b) for the W view, with similar distributions for the U and V views, which are not shown. The off-diagonal contribution where errors are made is typically found in a class adjacent to the true one. The region in which most errors are made is that in the immediate vicinity of the true interaction vertex, which is unsurprising given the relative infrequency of hits in this region (geometric considerations result in fewer pixels belonging to the classes closest to the vertex).
@@
The behaviour of the network also considered the distribution of weights in the trained model, the structure of the loss landscape in the vicinity of the minimum and we also explored the evolution of the learned weights in various convolutional layers to assess the stability of the training process. Fig. 5(c) shows the distribution of weights in the trained pass 1, W view network, with many small weights ensuring that the network can generalise to previously unseen events. Fig. 6 shows how the average loss over 1024 validation events varies as we take steps away from the minimum using the method described by Li et al [23]. The loss landscape (note the 3D surface depicts the logarithm of the average loss) appears smooth over the full grid of assessed weights and zooming into the region close to the minimum the landscape continues to exhibit features that lead to effective training.
@@ 6 Vertex reconstruction performance
In summary, the vertex reconstruction performance exhibits an evolution whereby performance is lowest for those events where there is little information in the vicinity of the vertex. This is particularly acute for NC interactions with low inelasticity and low hadronic invariant mass. The resultant lack of track-like deposition in the vicinity of the vertex means there is no handle with which to infer the vertex location. In CC interactions, these issues are largely offset by the presence of a leading lepton, whose charge deposition visibly leads back to the interaction vertex. As events increase in complexity, the increased particle multiplicity associated with events that have higher inelasticity or hadronic invariant mass leads to increases in vertex reconstruction performance. More charge deposition leading back to the vertex naturally makes the vertex location easier to identify, particularly with NC interactions, which start with the least available information. Eventually, however, increasing inelasticity or hadronic invariant mass increases the number of secondary and tertiary vertices, which act as additional candidate vertices, and leads to more overlapping particle trajectories that smear the paths back to the true interaction vertex, which can lead to a plateauing, and even a slight over-turning of performance at the highest inelasticities and hadronic invariant masses in CC interactions.
7 Model dependence
@@
Figure 12: The proton multiplicity for the different samples used for model dependency checks.
@@
Figure 14: 1.6 GeV NC interaction with a $\pi^{0}\rightarrow\gamma\gamma$, nine neutrons and either (left) three or (right) zero protons in the final state. The true interaction vertex is indicated by the blue circle (hidden below the reconstructed vertex in the left image), while the reconstructed interaction vertex is indicated by the red circle.
Figure 15: 2.8 GeV NC interaction with a $\pi^{0}\rightarrow\gamma\gamma$ and either (left) a neutron or (right) proton in the final state. The true interaction vertex is indicated by the blue circle, while the reconstructed interaction vertex is indicated by the red circle.
@@
Figure 18: 24.9 GeV CC interaction with a $\mu$, a $\pi^{+}$ and either (left) a neutron or (right) proton in the final state. The true interaction vertex is indicated by the blue circle, while the reconstructed interaction vertex is indicated by the red circle.
@@ 9 Conclusion
Support for deep neural networks has been integrated into the Pandora pattern recognition reconstruction workflow using LibTorch. A U-ResNet classifying hits with respect to their distance to the neutrino interaction vertex has been implemented in the context of a DUNE horizontal-drift far detector and acts as a performant vertex finding technique that substantially outperforms the previous BDT implementation, with an increase in the efficiency of vertex reconstruction within 1 cm of the true vertex of more than 20% in all flavours. Given the transferability of Pandora’s algorithms to other detector contexts, it is expected that this approach will also be effective in a vertical-drift far detector, though perhaps with a small reduction in resolution reflecting the larger induction channel spacing in the vertical-drift detector. Charged current interactions yield highly performant vertex finding, while neutral current performance is reduced by the absence of pointing information from a leading lepton (though still much improved relative to the BDT implementation). Interaction vertices are identified with equivalent efficiency in $\nu_{e}$ and $\nu_{\mu}$ samples, while $\nu_{\tau}$ performance is reduced due to a larger neutral current fraction. In general, precise vertex reconstruction for events with little charge deposition in the vicinity of the neutrino interaction vertex is very difficult, as expected, but vertex reconstruction performance rapidly improves as particle multiplicity in this region increases. This performance then plateaus and even over-turns for the most complex events with secondary and tertiary vertices acting as plausible alternative candidates, and overlapping particle trajectories smearing the path back to the true interaction vertex. Evaluation of model dependence in terms of final state proton multiplicity shows no direction bias in the reconstructed vertex position. Though neutral current interactions yield lower vertex efficiency (more catastrophic failures) as the proton multiplicity goes to zero (which is to be expected given reduced pointing information), charged current interactions appear insensitive to proton multiplicity. These improvements in vertex reconstruction will facilitate improvements in hit clustering, particle characterisation and subsequent high-level reconstruction quantities, such as estimates of neutrino energy, by avoiding inappropriate splitting and merging of particles, and errors in parent-child relationships.

### v2

@@ 1 Introduction
Figure 1: Schematic of the workspace geometry (red) within the context of the full detector geometry (black). APAs are indicated by gray boxes in the YZ plane of the workspace volume, while CPA walls are indicated in red in the YZ plane. APA walls in the full geometry are indicated by the black boxes in the YZ plane.
Section 2 briefly introduces LArTPCs, while Section 3 describes the simulated data used for this work. The conceptual approach to finding the vertex and the corresponding truth definition are presented in Section 4. Section 5 summarises the network architecture, provides details of the training metrics and assesses the behaviour of the loss function and network weights. Results in a simulated experimental environment are presented in Section 6. The performance characteristics, sources of bias and robustness considerations are discussed in Section 7 and a number of extensions/alternative approaches planned for future work are described in Section 8.
@@ 2 Liquid argon time projection chambers
Two far detector designs currently exist for DUNE, the horizontal-drift [7] and vertical-drift [8] detectors. These two designs utilise different geometries, orientations, drift lengths and readout technologies, the details of which are left to the aforementioned references. Conceptually the operation is the same, and Pandora is agnostic to the detector differences. A LArTPC volume is a fully active liquid argon target, with a uniform electric field applied between a cathode and anode. Charged particles propagating through this medium ionise the argon (see Fig. 2), producing drift electrons and scintillation photons. The drift electrons are carried by the electric field to the anode where they are read out on three readout planes: Two planes (U and V) with voltage biased to be transparent to the electrons and thereby allow current to be induced in readout channels as the electrons pass by, and a readout plane (W) with voltage set to permit collection of the drift electrons on the readout channels. The scintillation light is detected by photon detectors embedded in anode and cathode planes [9], though Pandora currently only reads charge information coming from the TPC waveforms, and does not yet read in scintillation light information and therefore this will not be discussed further in this article.
@@ 5 Training and understanding the network
Given the previously described formulation of the problem in terms of semantic segmentation, the architecture used has a dense U-ResNet structure [19, 21]. The network has four blocks in the down-sampling path (extracting features) and in the up-sampling path (recovering spatial detail), with all convolutional filters of size 3x3. The convolutional blocks implement residual skip connections between the input to those blocks and the final ReLU activation function of the respective block (adding the input to the block to the intermediate network output prior to the activation function). Due to the limited training batch size of 32 images per batch required to fit within GPU memory (16 GB) during training, group, rather than batch, normalisation is applied within convolutional and transpose convolutional blocks to ensure the inputs to network layers don’t drift as training proceeds.
@@
For the second pass training, with equivalent inclusion cuts applied, there were 85981, 85555, and 84658 for the U, V and W views, respectively. Note that images are excluded if errors in the first pass produce framing of the candidate vertex region that does not contain the true vertex. For validation there were 28660, 28518, and 28219 for the U, V and W views, respectively. Each network was trained for 25 epochs and the evolution of the network performance across each of those training iterations is shown in Fig. 5(a)-5(d).
@@
Also of interest are the confusion matrices for the various classes of the different networks. As already noted, per class accuracy is high, but one can also observe the distribution of classification errors in Fig. 6(a)-6(b) for the W view, with similar distributions for the U and V views, which are not shown. The off-diagonal contribution where errors are made is typically found in a class adjacent to the true one. The region in which most errors are made is that in the immediate vicinity of the true interaction vertex, which is unsurprising given the relative infrequency of hits in this region (geometric considerations result in fewer pixels belonging to the classes closest to the vertex).
@@
The behaviour of the network also considered the distribution of weights in the trained model, the structure of the loss landscape in the vicinity of the minimum and we also explored the evolution of the learned weights in various convolutional layers to assess the stability of the training process. Fig. 6(c) shows the distribution of weights in the trained pass 1, W view network – small weights reduce the sensitivity of the network to small perturbations [23], helping the network to generalise to previously unseen events. Fig. 7 shows how the average loss over 1024 validation events varies as we take steps away from the minimum using the method described by Li et al [24]. The loss landscape (note the 3D surface depicts the logarithm of the average loss) appears smooth over the full grid of assessed weights and zooming into the region close to the minimum the landscape continues to exhibit features that lead to effective training.
@@ 6 Vertex reconstruction performance
In summary, the vertex reconstruction performance exhibits an evolution whereby performance is lowest for those events where there is little information in the vicinity of the vertex, as expected. This is particularly acute for NC interactions with low inelasticity and low hadronic invariant mass, whereas these issues are largely offset in CC interactions by the presence of a leading lepton. As inelasticity or hadronic invariant mass increase, performance improves, as higher particle multiplicity yields more charge deposition leading back to the vertex, and naturally makes vertex identification easier. Eventually, however, increases in the number of secondary vertices, which act as additional candidates, and more overlapping particle trajectories smearing the paths back to the true interaction vertex, limit performance at the highest inelasticities and hadronic invariant masses in CC interactions.
7 Robustness testing
@@
Figure 13: The proton multiplicity for the different samples used for robustness checks.
@@
Figure 15: 1.6 GeV NC interaction with a $\pi^{0}\rightarrow\gamma\gamma$, nine neutrons and either (left) three or (right) zero protons in the final state. The true interaction vertex is indicated by the blue circle (hidden below the reconstructed vertex in the left image), while the reconstructed interaction vertex is indicated by the red circle. Particle colours are arbitrary and not correlated between left and right.
Figure 16: 2.8 GeV NC interaction with a $\pi^{0}\rightarrow\gamma\gamma$ and either (left) a neutron or (right) proton in the final state. The true interaction vertex is indicated by the blue circle, while the reconstructed interaction vertex is indicated by the red circle. Particle colours are arbitrary and not correlated between left and right.
@@
Figure 19: 24.9 GeV CC interaction with a $\mu$, a $\pi^{+}$ and either (left) a neutron or (right) proton in the final state. The true interaction vertex is indicated by the blue circle, while the reconstructed interaction vertex is indicated by the red circle. Particle colours are arbitrary.
@@ 9 Conclusion
Support for deep neural networks has been integrated into the Pandora pattern recognition reconstruction workflow using LibTorch. A U-ResNet classifying hits with respect to their distance to the neutrino interaction vertex has been implemented in the context of a DUNE horizontal-drift far detector and acts as a performant vertex finding technique that substantially outperforms the previous BDT implementation, with an increase in the efficiency of vertex reconstruction within 1 cm of the true vertex of more than 20% in all flavours. Given the transferability of Pandora’s algorithms to other detector contexts, it is expected that this approach will also be effective in a vertical-drift far detector, though perhaps with a small reduction in resolution reflecting the larger induction channel spacing in the vertical-drift detector. Charged current interactions yield highly performant vertex finding, while neutral current performance is reduced by the absence of pointing information from a leading lepton (though still much improved relative to the BDT implementation). Interaction vertices are identified with equivalent efficiency in $\nu_{e}$ and $\nu_{\mu}$ samples, while $\nu_{\tau}$ performance is reduced due to a larger neutral current fraction. In general, precise vertex reconstruction for events with little charge deposition in the vicinity of the neutrino interaction vertex is very difficult, as expected, but vertex reconstruction performance rapidly improves as particle multiplicity in this region increases. This performance then plateaus and even over-turns for the most complex events with secondary and tertiary vertices acting as plausible alternative candidates, and overlapping particle trajectories smearing the path back to the true interaction vertex. Evaluation of robustness in terms of final state proton multiplicity shows no direction bias in the reconstructed vertex position. Though neutral current interactions yield lower vertex efficiency (more catastrophic failures) as the proton multiplicity goes to zero (which is to be expected given reduced pointing information), charged current interactions appear insensitive to proton multiplicity. These improvements in vertex reconstruction will facilitate improvements in hit clustering, particle characterisation and subsequent high-level reconstruction quantities, such as estimates of neutrino energy, by avoiding inappropriate splitting and merging of particles, and errors in parent-child relationships.
Data availability statement
This manuscript has no associated data or the data will not be deposited. [Authors’ comment: The Pandora event reconstruction described in this paper operates on low-level data and does not directly produce physics results, hence there is no associated data release.]
Code availability statement
This manuscript has associated code/software in a data repository. [Authors’ comment: The Pandora event reconstruction software can be found at https://github.com/PandoraPFA.]